ADNAN MUHAMMED

Grade 12,

The plates of a capacitor of capacitance 10 μF, charged to 60 μC, are joined together by a wire of resistance 10 Ω at t = 0. Find the charge on the capacitor in the circuit at (a) t = 0, (b) t = 30 μs, (c) t = 120 μs and (d) t = 1.0 ms.

The plates of a capacitor of capacitance 10 μF, charged to 60 μC, are joined together by a wire of resistance 10 Ω at t = 0. Find the charge on the capacitor in the circuit at (a) t = 0, (b) t = 30 μs, (c) t = 120 μs and (d) t = 1.0 ms.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. C = 10 μF, Q = 60 μC, R = 10 Ω a) at t = 0, q = 60 μc b) at t = 30 μs, q = Qe^–t/RC = 60 * 10^–6 * e^–0.3 = 44 μc c) at t = 120 μs, q = 60 * 10^–6 * e^–1.2 = 18 μc d) at t = 1.0 ms, q = 60 * 10^–6 * e^–10 = 0.00272 = 0.003 μc.